OpenStudy (whimsical):

Help pls i cant solve this question: l 2-iz l = 2 i am supposed to draw the loci of z using that equation.

OpenStudy (anonymous):

Maybe this can help you. The locus of |2-z| =2 is a circle of radius 2 and center (2,0i). If you have a complex nummber z and you multiply it by another whoes modulus is 1 , you get another number whoes argument is the summ of the arguments. So zi is a nummber z turned 90º counterclockwise.

OpenStudy (whimsical):

the ans is circle with centre coordinates (0,-2) radius 2

OpenStudy (anonymous):

try this: if z= x+iy \[\sqrt{(2-y)^{2}+x ^{2}} = 2\]

OpenStudy (anonymous):



OpenStudy (whimsical):

ok i understood thank you for your guidance.

OpenStudy (anonymous):

can you "guide " me on my question now?

OpenStudy (anonymous):

2º from the top